### v1

# LETS-C: Leveraging Language Embedding for Time Series Classification

Abstract
Recent advancements in language modeling have shown promising results when applied to time series data. In particular, fine-tuning pre-trained large language models (LLMs) for time series classification tasks has achieved state-of-the-art (SOTA) performance on standard benchmarks. However, these LLM-based models have a significant drawback due to the large model size, with the number of trainable parameters in the millions. In this paper, we propose an alternative approach to leveraging the success of language modeling in the time series domain. Instead of fine-tuning LLMs, we utilize a language embedding model to embed time series and then pair the embeddings with a simple classification head composed of convolutional neural networks (CNN) and multilayer perceptron (MLP). We conducted extensive experiments on well-established time series classification benchmark datasets. We demonstrated LETS-C not only outperforms the current SOTA in classification accuracy but also offers a lightweight solution, using only 14.5% of the trainable parameters on average compared to the SOTA model. Our findings suggest that leveraging language encoders to embed time series data, combined with a simple yet effective classification head, offers a promising direction for achieving high-performance time series classification while maintaining a lightweight model architecture.
1 Introduction
Time series classification Bagnall et al. (2017); Abanda et al. (2019); Ismail Fawaz et al. (2019) has gained significant attention in recent years due to its wide-ranging applications in various domains, such as finance Passalis et al. (2017), healthcare Lipton et al. (2016), and activity recognition Yang et al. (2015). The increasing availability of time series data has driven the need for efficient and accurate classification methods. Recent advancements in natural language processing (NLP) and large language models (LLMs) have shown strong promises in language modeling Achiam et al. (2023), particularly in capturing temporal dependencies within sequential data. Inspired by this success, researchers have explored extending these techniques to the time series domain by fine-tuning pre-trained LLMs Zhou et al. (2024); Jin et al. (2023), achieving state-of-the-art (SOTA) performance on well-established benchmarks for tasks including classification and forecasting.
However, the use of LLMs for time series classification comes with a significant drawback due to their large model size. These models often have billions of parameters, making them computationally expensive and limiting their usage in resource-constrained environments Bommasani et al. (2021). At training time, fine-tuning partially frozen pre-trained LLMs also often involves millions of trainable parameters Zhou et al. (2024). To address this challenge, we propose an alternative approach to leverage the success of language modeling in the time series domain. In particular, we propose a novel approach called LETS-C (Language Embeddings for Time Series Classification) that leverages language embeddings, also known as text embeddings, for time series classification. To the best of our knowledge, this work is the first to explore the potential of language embeddings in time series analysis, specifically classification, and demonstrate SOTA performance.
LETS-C combines language embeddings with a simple yet effective classification head composed of convolutional neural networks (CNNs) and a multilayer perceptron (MLP). By projecting time series data using language embedding models, we capture the intricate patterns and dependencies present in the temporal data. The embeddings and time series are then fed into the classification head, which learns to discriminate between different classes. Through extensive experiments on a standard benchmark containing 10 datasets across various domains Bagnall et al. (2018), we demonstrated that LETS-C outperforms 20 baselines including the previous SOTA method. Moreover, LETS-C is significantly more efficient, using much less trainable parameters than the previous SOTA method.
Our main contributions are as follows:
  • Language Embeddings for Time Series: We introduce LETS-C, the first work to leverage language embeddings (or text embeddings) for time series analysis, specifically for classification tasks.
  • State-of-the-Art Performance: LETS-C achieves SOTA performance in classification accuracy on a well-established benchmark containing 10 datasets across different domains, surpassing 20 baselines.
  • Lightweight: LETS-C is significantly more efficient, achieving higher accuracy with much fewer trainable parameters (14.5%) compared to the existing SOTA method.
In addition, we conducted comprehensive analysis to along different aspects to showcase the effectiveness of LETS-C:
  • Text Embeddings Models: LETS-C with different text embedding models consistently outperforms previous SOTA with much fewer trainable parameters, further validating the effectiveness of our approach.
  • Time Series Embeddings: We revealed the advantage of text embeddings on time series, showing the embeddings of time series from the same class are more similar than the ones from different classes, explaining the boost in classification accuracy
  • Trade-off between Accuracy and Model Size: LETS-C retains a high percentage of accuracy even when the model size shrinks considerably, making it even more computationally efficient without compromising too much on model accuracy.
The remainder of this paper is organized as follows: Section 2 briefly summarizes related work. Section 3 details the structure of the proposed approach. In Section 4, we describe the experimental setup, benchmark results, and additional findings from comprehensive evaluations. Finally, Section 5 discusses our findings and Section 6 outlines limitations and future research directions. Due to space limitations, a more extensive discussion of the experimental setup and results is provided in the Appendix.
2 Related Work
In this section, we review the related works in three key areas: time series classification, the application of language models to time series data, and text embeddings.
Time Series Classification Time series classification has been an active research area for decades. We include an extended review in Appendix A due to space limit. Early methods focused on distance-based approaches Abanda et al. (2019), such as Dynamic Time Warping (DTW) Berndt & Clifford (1994) and distance kernels with Support Vector Machines (SVMs) Kampouraki et al. (2008). Others extracted features and used linear or tree-based classifiers like eXtreme Gradient Boosting (XGBoost) Chen & Guestrin (2016). Later, deep learning-based approaches, including Convolutional Neural Networks (CNNs) Wu et al. (2022a); Zhao et al. (2017), Multilayer Perceptron (MLP) Zhang et al. (2022), and Recurrent Neural Networks (RNNs) like Long Short-Term Memory (LSTM) Lai et al. (2018), gained popularity. These models can learn complex patterns and handle long sequences. Recently, Transformer-based models Vaswani et al. (2017) have been adapted from Natural Language Processing (NLP) to the time series domain Zhou et al. (2022); Wu et al. (2021); Zhou et al. (2021). They use self-attention mechanisms to model long-range dependencies. However, these more complex models often have larger sizes and higher computational costs, especially for training.
Language Models for Time Series The success of language modeling in NLP and LLMs has inspired researchers to harness LLMs in the time series domain. Comprehensive surveys Zhang et al. (2024); Jiang et al. (2024) have offered valuable insights into the integration of LLMs in time series analysis, highlighting key methodologies, challenges, and future directions.  Gruver et al. (2023); Xue & Salim (2023) enabled pre-trained LLMs to generate time series forecasts through prompting. Further,  Yu et al. (2023) delved into the potential of LLMs for generating explainable forecasts of financial time series.  Jin et al. (2023) introduced Time-LLM, which focuses on learning to project time series into the language embedding space and directly using pre-trained LLM for time series forecasting tasks. More importantly, recent work OneFitsAll by Zhou et al. (2024) has shown promising results by fine-tuning models like GPT Radford et al. (2019) on time series tasks. OneFitsAll achieved SOTA performance compared across various time series tasks including classification. Inspired by these successes, we propose to explore another form of language modeling, text embeddings, in the time series domain. We discover that this direction leads to new SOTA performance on time series classification tasks, with the benefit of much fewer trainable parameters than OneFitsAll.
Text Embeddings Text embeddings have played a crucial role in NLP. These embeddings map words or sentences into a dense vector space, capturing semantic and syntactic information. Various text embedding techniques have been proposed, ranging from word-level embeddings like Word2Vec Mikolov et al. (2013) and GloVe Pennington et al. (2014) to contextualized embeddings obtained from pre-trained language models such as BERT Devlin et al. (2018) and RoBERTa Liu et al. (2019b). In the time series domain, some works have proposed unsupervised methods for learning time series embeddings Franceschi et al. (2019); Tonekaboni et al. (2021). However, the availability of large-scale datasets in the time series domain is generally more limited compared to those in the NLP domain, making learning time series embedding from scratch more challenging compared to text embeddings. To our best knowledge, we are the first to leverage the well-trained text embeddings from the NLP domain for time series classification.
3 Methodology
Given a time series classification dataset $\mathcal{D}=\{(\mathbf{x_{i}},y_{i})_{i=1}^{N}\}$, where $\mathbf{x_{i}}$ is a multivariate time series sample, and $y_{i}\in\{1,2,\dots,C\}$ is the corresponding class label, the goal is to learn a classifier that accurately predicts the class label $\hat{y}_{i}$ for each time series. As illustrated in Figure 1, we propose LETS-C framework that harnesses text embeddings for time series classification tasks. Specifically, we 1) initially preprocess the time series data to standardize it, then 2) subsequently generate text embeddings from the standardized time series, 3) fuse embeddings with the time series data, and finally 4) feed the fused representation to a classification head that consists of CNNs and MLP. The choice of a simple classification head is intentional, we aim to test the hypothesis that the text embeddings of the time series provide sufficiently powerful representations for effective classification.
Figure 1: Left: Conventional text embedding. Right: Our proposed LETS-C framework normalizes time series and formats them to tokenize each digit separately. It then embeds the time series into the embedding space, fuses the embeddings and time series together via element-wise addition, and employs a simple classification head consisting of a CNN and an MLP for classification. In this framework, the only elements trained are the lightweight CNN model and the MLP head.
Preprocessing To ensure consistent scales across all model inputs, each feature dimension of time series $\mathbf{x_{i}}$ is min-max normalized to the range $[0,1]$ based on the minimum and maximum feature values of each dimension across the training data.
Text Embedding of Time Series It is crucial to carefully format the preprocessed time series into strings before using text embeddings, as the tokenization of numerical strings can significantly affect the embeddings.  Liu & Low (2023) has shown that tokenization impacts a model’s arithmetic abilities, with commonly used subword tokenization methods like Byte Pair Encoding (BPE) arbitrarily subdividing numbers, causing similar numbers to appear very differently. To mitigate this, we adopted a digit-space tokenization strategy, as suggested by Gruver et al. (2024), where each digit is spaced, commas are added to separate time steps, and decimal points are omitted for fixed precision. For instance, a series with a precision of two decimal places, such as 0.645, 6.45, 64.5, 645.0, would be formatted as "6 4 , 6 4 5 , 6 4 5 0 , 6 4 5 0 0" prior to tokenization. This method ensures separate tokenization of each digit, preserving numerical integrity and enhancing pattern recognition in language models.
Next, we utilized OpenAI’s text-embedding-3-large model OpenAI (2024) to embed the formatted time series into the embedding space. This model was selected for several reasons: it is highly ranked on the Massive Text Embedding Benchmark (MTEB) leaderboard MTEB (2024); Muennighoff et al. (2022), known for its effectiveness in a variety of downstream tasks such as text search and sentence similarity; it supports a high maximum token length of 8191, accommodating our time series datasets; and it offers a high-dimensional vector space of 3072 dimensions. This large dimensionality captures a broad spectrum of temporal features, yet the model allows for truncation to reduce dimensions as needed for specific applications without substantial loss of semantic information Kusupati et al. (2022). This capability to truncate dimensions is particularly advantageous for optimizing computational efficiency while maintaining robust performance in downstream applications, aligning well with our goal of a lightweight framework.
In particular, we take each dimension of $\mathbf{x_{i}}$ and generate the corresponding text embedding. Thus, we transform $\mathbf{x_{i}}\in\mathbb{R}^{d\times l_{x}}$ into embeddings $\mathbf{e_{i}}\in\mathbb{R}^{d\times l_{e}}$, where $d$ is the multivariate dimension, $l_{x}$ is the length of the time series, and $l_{e}$ is the length of the embedding. Alternatively, each time series can be divided into separate patches where each patch has a corresponding text embedding. However, this alternative is not scalable and did not yield significant performance benefits empirically. Note that the embedding computation in LETS-C is a one-time pass, contrasting with the persistent computational cost by models such as OneFitsAll which rely on partially frozen transformer-based components.
To validate that text embeddings are suitable for time series classification, we compared the cosine similarities of text embeddings of time series from the same and different classes. Our results show consistently higher intra-class similarity than inter-class. This highlights the effectiveness of text embeddings in time series tasks (see Section 4.3).
Fusing Embedding and Time Series Next, we perform an element-wise addition of the embedding to the preprocessed time series, while preserving the maximum length between the two. Specifically, $\mathbf{x_{i}}\in\mathbb{R}^{d\times l_{x}}$ are added with embeddings $\mathbf{e_{i}}\in\mathbb{R}^{d\times l_{e}}$. If $l_{x}$ and $l_{e}$ do not match, we pad the shorter one with zeros to align their sizes before addition, resulting in a combined representation in $\mathbb{R}^{d\times max(l_{x},l_{e})}$. The direct addition of embeddings to the time series data is analogous to the use of positional embeddings in NLP, which enriches the sequence representation with additional structural and contextual information. As a result, it enables both time series and text embeddings to contribute to the learning process, thereby enhancing the model’s ability to effectively interpret complex temporal patterns (Section 4.3 show alternative fusion approaches are less effective).
Lightweight Classification Head Lastly, we pair the fused embedding and time series with a simple classification head composed of 1D CNNs and an MLP for time series classification. The output from CNNs are flattened and fed through the final MLP head, which outputs a vector of the probabilities of each time series class. Hyperparameter search determines the number of convolutional blocks in the CNN, the number of linear layers in the MLP, and the utilization of batch normalization, dropout, activation functions, and pooling operations. With a simple classification head, our model is lightweight and requires much less trainable parameters compared to the existing SOTA method built on transformers, as benchmarked in Section 4.2.
4 Experiments
We provide a brief introduction to the datasets, baselines, evaluation metrics, and model implementation details in Section 4.1. In Section 4.2, we compare our approach against various baselines—including classical methods and MLP-, CNN-, RNN-, and Transformer-based techniques, as well as recent SOTA LLM-based methods—highlighting LETS-C’s performance and computational costs. Further insights into ablation studies of LETS-C, generalization of our method across different embedding types, the power of the embeddings and its similarities across different time series are explored through additional analyses in Section 4.3.
4.1 Experimental Setup
Datasets and Evaluation Metrics We followed Wu et al. (2022a) and benchmarked on 10 multivariate datasets from the UEA Time Series Classification Archive (Bagnall et al., 2018), encompassing a wide range of applications, namely: EthanolConcentration (Large et al., 2018), FaceDetection (Rik Henson, 2023), Handwriting (Shokoohi-Yekta et al., 2017), Heartbeat (Liu et al., 2016), JapaneseVowels (Kudo et al., ), PEMS-SF (Cuturi, 2011), SelfRegulationSCP1 (Birbaumer et al., 1999), SelfRegulationSCP2 (Birbaumer et al., 1999), SpokenArabicDigits (Bedda & Hammami, 2010), and UWaveGestureLibrary (Liu et al., 2009). See Table 6 for their data characteristics. These datasets provide a comprehensive testing environment, featuring multivariate dimensions ranging from 3 to 963, time series lengths up to 1751, and number of classes up to 26. See Section B in the Appendix for details on each dataset.
To benchmark the classifiers, we used metrics including classification accuracy and AvgWins. AvgWins is defined as the average number of times that a method outperforms other methods across benchmarked datasets, with ties also being counted towards this average. Additionally, the methods were analyzed in terms of trainable model parameters.
Baselines We included 20 baseline models to ensure a comprehensive comparison. We utilize the same baselines outlined by Zhou et al. (2024); Wu et al. (2022a), namely: Classical methods: 1) Dynamic Time Warping (DTW) (Berndt & Clifford, 1994), 2) eXtreme Gradient Boosting (XGBoost) (Chen & Guestrin, 2016), and 3) RandOm Convolutional KErnel Transform (ROCKET) (Dempster et al., 2020); MLP-based methods: 4) LightTS (Zhang et al., 2022), and 5) DLinear (Zeng et al., 2023); RNN-based models: 6) Long short-term memory (LSTM) (Hochreiter & Schmidhuber, 1997), 7) LSTNet (Lai et al., 2018), and 8) LSSL (Gu et al., 2021); CNN-based models: 9) Temporal Convolutional Network (TCN) (Franceschi et al., 2019), and 10) TimesNet (Wu et al., 2022a); Transformer-based models: 11) Transformer (Vaswani et al., 2017), 12) Reformer (Kitaev et al., 2020), 13) Informer (Zhou et al., 2021), 14) Pyraformer (Liu et al., 2021), 15) Autoformer (Wu et al., 2021), 16) Stationformer (Liu et al., 2022), 17) FEDformer (Zhou et al., 2022), 18) ETSformer (Woo et al., 2022), 19) Flowformer (Wu et al., 2022b); and LLM-based model: 20) OneFitsAll (Zhou et al., 2024). Note that some of these methods were originally developed for forecasting but were later adapted for classification tasks.
Implementation Details The experiments were conducted on a Linux machine equipped with 8 NVIDIA T4 Tensor Core GPUs, each with 16GB of memory. We utilized the PyTorch v1.0.0 deep learning platform running on Python 3.6 for all models. All configurations employed the RAdam optimizer Liu et al. (2019a), with its default hyperparameters settings $(\beta_{1},\beta_{2})=(0.9,0.999)$. For tokenization, we found that maintaining a precision of one decimal place optimizes performance (see Appendix Section F). Further, we explored LETS-C’s performance across various embedding types, other than text-embedding-3-large in Section 4.3. Exploratory hyperparameter optimization was conducted, revealing that 1-4 1D convolutional layers and 1-3 linear layers are optimal for the performance of LETS-C across all datasets.
4.2 Main Results
Comparison to State-of-the-art Table 1 presents a comparative analysis of our proposed LETS-C approach against 20 baseline models across 10 benchmark datasets as introduced above.
Table 1: Comparison of classification accuracy (%) and AvgWins (%). Red: Best, Blue: Second best.
[…]
Classical methods DTW Berndt & Clifford (1994) 32.3 52.9 28.6 71.7 94.9 71.1 77.7 53.9 96.3 90.3 66.97 0%
XGBoost Chen & Guestrin (2016) 43.7 63.3 15.8 73.2 86.5 98.3 84.6 48.9 69.6 75.9 65.98 10%
ROCKET Dempster et al. (2020) 45.2 64.7 58.8 75.6 96.2 75.1 90.8 53.3 71.2 94.4 72.53 20%
MLP LightTS Zhang et al. (2022) 29.7 67.5 26.1 75.1 96.2 88.4 89.8 51.1 100 80.3 70.42 10%
DLinear Zeng et al. (2023) 32.6 68 27 75.1 96.2 75.1 87.3 50.5 81.4 82.1 67.53 0%
RNN LSTM Hochreiter & Schmidhuber (1997) 32.3 57.7 15.2 72.2 79.7 39.9 68.9 46.6 31.9 41.2 48.56 0%
LSTNet Lai et al. (2018) 39.9 65.7 25.8 77.1 98.1 86.7 84 52.8 100 87.8 71.79 10%
LSSL Gu et al. (2021) 31.1 66.7 24.6 72.7 98.4 86.1 90.8 52.2 100 85.9 70.85 10%
CNN TCN Franceschi et al. (2019) 28.9 52.8 53.3 75.6 98.9 68.8 84.6 55.6 95.6 88.4 70.25 0%
TimesNet Wu et al. (2022a) 35.7 68.6 32.1 78 98.4 89.6 91.8 57.2 99 85.3 73.57 0%
Transformers Transformer Vaswani et al. (2017) 32.7 67.3 32 76.1 98.7 82.1 92.2 53.9 98.4 85.6 71.9 0%
Reformer Kitaev et al. (2020) 31.9 68.6 27.4 77.1 97.8 82.7 90.4 56.7 97 85.6 71.52 0%
Informer Zhou et al. (2021) 31.6 67 32.8 80.5 98.9 81.5 90.1 53.3 100 85.6 72.13 20%
Pyraformer Liu et al. (2021) 30.8 65.7 29.4 75.6 98.4 83.2 88.1 53.3 99.6 83.4 70.75 0%
Autoformer Wu et al. (2021) 31.6 68.4 36.7 74.6 96.2 82.7 84 50.6 100 85.9 71.07 10%
Stationformer Liu et al. (2022) 32.7 68 31.6 73.7 99.2 87.3 89.4 57.2 100 87.5 72.66 20%
FEDformer Zhou et al. (2022) 31.2 66 28 73.7 98.4 80.9 88.7 54.4 100 85.3 70.66 10%
ETSformer Woo et al. (2022) 28.1 66.3 32.5 71.2 95.9 86 89.6 55 100 85 70.96 10%
Flowformer Wu et al. (2022b) 33.8 67.6 33.8 77.6 98.9 83.8 92.5 56.1 98.8 86.6 72.95 0%
LLM OneFitsAll Zhou et al. (2024) 34.2 69.2 32.7 77.2 98.6 87.9 93.2 59.4 99.2 88.1 73.97 20%
LETS-C (Ours) 52.9 68.9 23.8 78 99.2 93.1 93.2 62.8 99.2 90.6 76.17 40%
We observe that LETS-C consistently demonstrates robust performance across all datasets, achieving the highest average accuracy of 76.16% and AvgWins of 40%, compared to 20 benchmark models. This includes the most recent SOTA model OneFitsAll (accuracy: 73.97%, AvgWins: 20%) and an older SOTA TimesNet (accuracy: 73.57%, AvgWins: 0%). Notably, LETS-C surpasses OneFitsAll on six out of ten datasets by a significant margin. LETS-C is particularly effective on challenging datasets like PEMS-SF and EthanolConcentration (EC). PEMS-SF is characterized by exceptionally high dimensionality with 963 features, and EC contains an extremely long time series at length of 1751, compared to other datasets in the benchmark. These results showcase the competitive edge of LETS-C against the previous SOTA methods (OneFitsALL), thus establishing a new benchmark for time series classification.
Computational Cost Analysis Next, we aim to assess how well the LETS-C approach balances performance with computational efficiency, which is crucial for usage in resource-constrained environments. Table 2 provides a detailed analysis of the trainable parameters associated with LETS-C compared to the previous SOTA model, OneFitsAll.
Table 2: Comparison of per batch trainable parameters (millions) for LETS-C vs. OneFitsAll. The Delta (%) $=100\times\frac{\text{Cost with LETS-C{}}}{\text{Cost with OneFitsAll}}$, quantifying the efficiency of LETS-C relative to OneFitsAll.
[…]
Trainable parameters (M) LETS-C (Ours) 0.28 0.003 0.15 0.04 0.14 0.56 0.30 0.33 0.14 0.26 0.22
OneFitsAll 1.42 2.37 1.73 2.03 1.32 10.23 0.98 1.04 1.82 1.0 2.39
Delta (%) 19.89 0.16 8.89 2.28 11.19 5.51 30.83 32.06 7.77 26.21 14.48
Our method achieved higher performance with only 14.48% of the trainable model parameters on average, compared to OneFitsAll. Despite the advantage of OneFitsAll over other leading models like TimesNet and FEDformer on its reduced parameter count, OneFitsAll still requires much more computational cost than our approach, especially for training. Further experiments on training and inference times are detailed in Table 7 in the Appendix. We show that LETS-C offers a lightweight approach to the time series classification task while achieving the SOTA performance.
4.3 Additional Analysis
Ablation Study To empirically assess the advantages of fusing both text embeddings and time series data, as opposed to variants that only leverage either the text embedding or time series itself, we conducted an ablation study.
Table 3: Comparison of classification accuracy (%) for different configurations: ours (both embeddings and time series), embeddings only, and time series only. Red: Best.
[…]
LETS-C (Ours) 52.9 68.9 23.8 78 99.2 93.1 93.2 62.8 99.2 90.6 76.17 60%
   embedding only 38 59.4 20.1 80 99.2 92.5 88.7 63.9 99.2 88.4 72.94 40%
   time series only 42.6 69.6 25.1 76.1 98.1 89 93.2 58.3 98.9 83.8 73.47 30%
As shown in Table 3, we observe that the combination of both embeddings and time series achieves the highest average accuracy, at 76.17%, and the greatest number of AvgWins, compared to the ablated versions. These demonstrate the substantial performance gains from fusing embeddings with time series data, essential for optimal model accuracy.
Alternative Methods for Fusing Time Series with Embeddings We explored two additional methods for fusing time series and embeddings beyond simple addition. The first method involves a Fusion network that processes both embeddings and time series data through convolutional and dense layers, subsequently merging the features from both sources into a final dense network. The second method employs Concatenation, where the time series and embeddings are concatenated and processed through a lightweight classification head. Table 4 presents the classification accuracy and trainable model parameters for these variations.
Table 4: Comparison of classification accuracy (%) and trainable model parameters (millions) for alternative methods for fusing time series with embeddings. Higher AvgWins and averages signify superior performance, while lower averages suggest greater computational efficiency. Red: Best performance.
[…]
LETS-C (Addition) 52.9 68.9 23.8 78 99.2 93.1 93.2 62.8 99.2 90.6 76.17 70%
Fusion network 44.1 66.5 23.6 76.6 98.1 86.1 92.8 56.1 99.2 90.9 73.40 20%
Concatenation 43 65.1 22.5 79 98.9 93.1 93.9 58.9 99 88.8 74.22 30%
[…]
LETS-C (Addition) 0.28 0.003 0.15 0.04 0.14 0.56 0.30 0.33 0.14 0.26 0.22
Fusion Network 0.19 0.35 0.33 0.28 0.16 5.54 0.37 0.27 0.23 0.04 0.78
Concatenation 0.42 0.009 0.26 0.17 0.11 0.28 0.23 0.39 0.09 0.46 0.24
We observe that the addition approach in the LETS-C architecture achieves the highest average classification accuracy (76.11%) compared to the fusion network (73.40%) and concatenation (74.22%). It also records the best AvgWins at 70%. Further, the number of parameters increases with both the concatenation and fusion network approaches, resulting in additional complexity compared to the addition method. As our goal was to develop a lightweight model, we opted for the addition approach due to its simpler parameter structure.
Assessing Text Embeddings with Cosine Similarity To analyze text embeddings extracted from time series data and assess their effectiveness in distinguishing between and within data classes, we calculate the average cosine similarities for all pairs of time series within each class and similarly for between classes. Additionally, we average similarities from multiple features to effectively handle multivariate time series. This approach helps us quantify how closely time series in the embedding space are related, both within individual classes and between distinct groups. Figure 2 (left) visualizes these relationships through heatmaps. Each matrix entry is scaled using min-max normalization to range from 0 to 1, where warmer colors in the heatmap represent higher similarities and darker shades indicate lower similarities. Diagonal entries show within-class similarities, highlighting intra-class cohesion, while off-diagonal entries reveal between-class relationships.
Figure 2: Left: Heatmaps illustrating within-class and between-class cosine similarities of text embeddings derived from the training time series data in JapaneseVowels (far left) with 9 classes and SpokenArabicDigits (middle) with 10 classes. Right: Trade-off between the percentage of accuracy retention and model parameter retention relative to LETS-C’s optimal values across all datasets. The optimal LETS-C accuracy (%) and parameters (millions) for each dataset are detailed in the legend.
We observe that within-class similarity consistently exceeds across-class similarity, thereby validating the hypothesis that text embeddings effectively retain and convey significant information from the underlying time series data. This confirms their utility in discerning inherent patterns within the data.
Embedding Type To explore the generalization capabilities of our approach with various embeddings beyond OpenAI’s text-embedding-3-large, we evaluated the performance of LETS-C using three distinct embedding types. These were selected based on their high rankings in the MTEB Muennighoff et al. (2022); MTEB (2024) overall leaderboard performance, varying in embedding dimensions, maximum token lengths, and model sizes. The selected embeddings and their characteristics, along with text-embedding-3-large, are detailed in Table 9, and Section G in the Appendix. Table 5 presents detailed performance metrics (accuracy) and computational cost (trainable parameters) for these various embedding models in the LETS-C framework.
We observe that LETS-C consistently outperforms current baselines in terms of average classification accuracy and AvgWins, while utilizing only a fraction of the trainable model parameters across all explored embedding types, which vary in dimensionality and model size. Specifically, text-embedding-3-large achieves an average accuracy of 76.17%, while e5-mistral-7b-instruct, gte-large-en-v1.5, and nomic-embed-text-v1 achieve accuracies of 74.89%, 76.02%, and 75.12% respectively. These figures not only surpass the previous SOTA accuracy of 73.97% but also require significantly fewer trainable model parameters—just 14.48%, 15.62%, 9.24%, and 5.31% compared to OneFitsAll. Although preliminary, we observe some patterns between the computational cost of LETS-C with an embedding model and the embedding dimensionality: e5-mistral-7b-instruct requires the highest average trainable parameters (0.25M) for the largest embedding dimensionality of 4096. This is followed by text-embedding-3-large, which has an embedding dimensionality of 3072 and needs 0.22M trainable parameters, then gte-large-en-v1.5 with a dimensionality of 1024 requiring 0.19M, and lastly nomic-embed-text-v1 with 768 dimensions needing 0.09M trainable parameters. Consequently, our approach generalizes across diverse embedding types, demonstrates superior performance, and offers a more lightweight solution, outperforming baselines across a broad range of language embedding models. Additionally, we examined the impact of time series precision on tokenization and its effect on LETS-C’s performance, detailed in Appendix Section F.
Table 5: Comparison of LETS-C using various embedding models against OneFitsAll. Performances that surpass OneFitsAll are in bold with the best in Red. AvgWins scores above 50% denote consistent superiority, calculated as 1 for outperformance relative to OneFitsAll and 0 otherwise. Delta (%) $=100\times\frac{\text{Cost with each embedding model}}{\text{Cost with OneFitsAll}}$ measures computational efficiency relative to OneFitsAll.
[…]
OneFitsAll 34.2 69.2 32.7 77.2 98.6 87.9 93.2 59.4 99.2 88.1 73.97 --
LETS-C text-embedding-3-large 52.9 68.9 23.8 78 99.2 93.1 93.2 62.8 99.2 90.6 76.17 80%
e5-mistral-7b-instruct 55.5 68.7 23.3 77.6 99.2 84.4 93.9 59.4 98.5 88.4 74.89 60%
gte-large-en-v1.5 57.8 68.8 24.7 77.6 98.4 91.3 94.2 60 99 88.4 76.02 60%
nomic-embed-text-v1 52.9 68 24.8 76.6 99.2 88.4 93.9 59.4 98.6 89.4 75.12 60%
Computational Cost (Trainable Parameters (M)) $\downarrow$
[…]
OneFitsAll 1.42 2.37 1.73 2.03 1.32 10.23 0.98 1.04 1.82 1.0 2.39
LETS-C text-embedding-3-large 0.28 0.003 0.15 0.04 0.14 0.56 0.30 0.33 0.14 0.26 0.22
           Delta % 19.89 0.16 8.89 2.28 11.19 5.51 30.83 32.06 7.77 26.21 14.48
e5-mistral-7b-instruct 0.13 0.40 0.39 0.17 0.16 0.30 0.24 0.24 0.33 0.16 0.25
           Delta % 9.48 16.93 22.78 8.54 12.55 2.97 25.06 23.59 18.39 15.93 15.62
gte-large-en-v1.5 0.18 0.31 0.31 0.12 0.04 0.56 0.03 0.07 0.22 0.09 0.19
           Delta % 12.91 13.44 18.27 6.11 3.36 5.51 3.77 7.65 12.34 9.00 9.24
nomic-embed-text-v1 0.03 0.03 0.36 0.07 0.05 0.19 0.02 0.10 0.06 0.02 0.09
           Delta % 2.21 1.31 20.99 3.58 3.99 1.85 3.02 10.03 3.34 2.78 5.31
Trade-offs: Model Accuracy vs. Parameter Complexity Next, we aim to examine the trade-offs between model accuracy and parameter complexity in our approach. We explore the relationship between retained accuracy and parameter reduction across all datasets, aiming to demonstrate minimal loss in accuracy despite significant reductions in model parameters. Figure 2 (right) illustrates the trade-off between the percentage of accuracy retention and the percentage of model parameter retention, as compared to the optimal accuracy and parameters of our LETS-C model across all datasets (see Tables 1, and 7). For more detailed results, see Table 10 in the Appendix. This comparison quantifies the accuracy and computational efficiency of the reduced model relative to our best-performing model.
Figure 2 (right) underscores the efficiency of our LETS-C approach and demonstrates its effectiveness in retaining high accuracy with significantly reduced parameters across all datasets. The trade-off between model accuracy and parameter complexity is data-dependent, but generally, a substantial reduction in parameters leads to only a slight decrease in accuracy. We closely explore this trade-off in three specific datasets, namely, Heartbeat, PEMS-SF, and SpokenArabicDigits, to understand how these trade-offs manifest in different contexts, as detailed in Appendix Section H.
These examples illustrate that while fewer parameters generally lead to lower accuracy, the decrement is manageable, making LETS-C highly suitable for deployment in resource-constrained environments or applications requiring rapid processing with minimal computational overhead.
5 Conclusion
In this work, we introduced LETS-C, a novel approach that leverages language embeddings for time series classification. To the best of our knowledge, this is the first work to explore the potential of off-the-shelf text embeddings in time series analysis, specifically for classification tasks. By projecting time series data using text embedding models and utilizing a simple yet effective classification head, LETS-C achieves SOTA performance on a well-established benchmark containing 10 datasets across different domains, surpassing 20 baseline methods. Moreover, LETS-C is significantly more lightweight compared to the previous SOTA method, achieving higher accuracy while using much less trainable parameters. Through comprehensive analysis, we demonstrated the effectiveness of LETS-C along different aspects, including its robustness across various text embedding models, the advantage of text embeddings for time series classification, and the trade-off between accuracy and model size. We believe that the findings and insights from this work will inspire further exploration of language embeddings and their potential applications in time series domain, paving the way for the development of more powerful and efficient methods for various time series tasks.
6 Limitations and Future Work
In this work, we adopted the tokenization method proposed by Gruver et al. (2023); however, there may be tokenization strategies better suited specifically for time series data. Future research could investigate these alternatives to optimize performance further (Rajaraman et al., 2024). Additionally, while this study concentrates on using text embeddings exclusively for time series classification, to more comprehensively assess their universality, future research should explore the application of these techniques to a broader spectrum of time series analysis tasks, including forecasting, anomaly detection, and imputation. We expect the societal impact of our work to improve decision-making processes in various sectors, but also practitioners need be careful particularly in sensitive areas like healthcare and personal finance, which raises significant privacy issues. If data is not handled with strict privacy controls, there could be risks of unauthorized access and misuse of personal data.
Acknowledgments
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co and its affiliates (“J.P. Morgan”) and is not a product of the Research Department of J.P. Morgan. J.P. Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Appendix
[…]
Appendix A Related Works
Time series classification has been an active research area for decades. Early approaches investigated distance-based approaches Abanda et al. (2019) for time series classification. Some built nearest neighbor classifiers based on explicit time series distance measures such as Dynamic Time Warping (DTW) Wang et al. (2013); Jeong et al. (2011); Berndt & Clifford (1994). Others have used distance kernels instead and learned Support Vector Machines (SVMs) Kampouraki et al. (2008); Bahlmann et al. (2002); Shimodaira et al. (2001), or extracted features and learned linear Dempster et al. (2020) or tree-based classifiers such as eXtreme Gradient Boosting (XGBoost) Chen & Guestrin (2016).
Later, deep learning-based approaches are widely adopted because of their ability to learn complex patterns. Convolutional Neural Networks (CNNs) have proven to be successful in learning local patterns in time series data Wu et al. (2022a); Franceschi et al. (2019); Zhao et al. (2017). Similarly, Multilayer Perceptron (MLP) can provide simple but effective time series classifiers Zhang et al. (2022). Recurrent Neural Networks (RNNs), such as Long Short-Term Memory (LSTM) effectively handle long sequence modeling Gu et al. (2021); Lai et al. (2018); Hochreiter & Schmidhuber (1997).
More recently, Transformer-based models Vaswani et al. (2017) have revolutionized the NLP domain, and these models have been adapted to the time series domain Zhou et al. (2022); Wu et al. (2021); Zhou et al. (2021). The self-attention mechanism in transformers is known for modeling long-range dependencies in sequence data. However, the increasing complexity of these models often comes with larger model sizes and higher computational costs, especially for training.
Appendix B Datasets
We benchmark our model using the following 10 multivariate datasets from the UEA Time Series Classification Archive Bagnall et al. (2018). See Table 6 for their data characteristics.
Table 6: Dataset Characteristics. Abbreviations: EC: EthanolConcentration, FD: FaceDetection, HW: Handwriting, HB: Heartbeat, JV: JapaneseVowels, PEMS: PEMS-SF, SCP1: SelfRegulationSCP1, SCP2: SelfRegulationSCP2, SAD: SpokenArabicDigits, UW: UWaveGestureLibrary
Characteristic EC FD HW HB JV PEMS-SF SCP1 SCP2 SAD UW
Train Size 261 5890 150 204 270 267 268 200 6599 120
Test Size 263 3524 850 205 370 173 293 180 2199 320
Number of Dimensions 3 144 3 61 12 963 6 7 13 3
Series Length 1751 62 152 405 29 144 896 1152 93 315
Number of Classes 4 2 26 2 9 7 2 2 10 8
Type Spectro EEG Motion/Human Activity Recognition Audio Audio Occupancy rate EEG EEG Speech EEG
B.1 EthanolConcentration
EthanolConcentration (Large et al., 2018) comprises raw spectra from water-and-ethanol solutions contained within 44 unique, real whisky bottles, featuring ethanol concentrations of 35%, 38%, 40%, and 45%. Scotch Whisky regulations require a minimum alcohol content of 40%, a standard that producers adhere to in order to comply with labeling specifications. The dataset presents a classification task to identify the ethanol concentration from spectral readings of any given bottle. Each record includes three spectral readings from the same bottle and batch, obtained by positioning the bottle between a light source and a spectroscope. These spectral readings, which cover wavelengths from 226nm to 1101.5nm at a 0.5nm resolution, were recorded over a one-second integration time using a StellarNet BLACKComet-SR spectrometer. The methodology deliberately avoids optimizing for clarity or consistency in the spectral path, aiming to simulate the varied conditions typical of rapid screening tests that may be performed on batches of spirits for quality assurance.
B.2 FaceDetection
The FaceDetection dataset originates from a 2014 Kaggle competition (Rik Henson, 2023). The challenge involves identifying whether a subject is viewing a picture of a face or a scrambled image using magnetoencephalography (MEG) data, independent of the individual subject. This dataset specifically includes only the training portion from the competition, organized by patient. It comprises data from 10 training subjects (subject01 to subject10) and 6 testing subjects (subject11 to subject16). Each subject has approximately 580 to 590 trials, resulting in a total of 5,890 training trials and 3,524 test trials. Each trial features 1.5 seconds of MEG data, initiated 0.5 seconds before the stimulus is presented, and is associated with a class label—Face (class 1) or Scramble (class 0). The data were down-sampled to 250Hz and subjected to a high-pass filter at 1Hz, producing 62 observations per channel.
B.3 Handwriting
The Handwriting dataset (Shokoohi-Yekta et al., 2017) consists of motion data captured from a smartwatch while subjects wrote the 26 letters of the alphabet. Developed at the University of California, Riverside (UCR), this dataset includes 150 training cases and 850 test cases. It features six dimensions, comprising three accelerometer readings and three gyroscope readings.
B.4 Heartbeat
The Heartbeat dataset originates from the PhysioNet/CinC Challenge 2016 Liu et al. (2016) and consists of cardiac sound recordings from a diverse pool of participants, both healthy individuals and patients with cardiac conditions. Recordings were made in various settings, clinical and non-clinical, and captured from multiple body locations including the aortic, pulmonic, tricuspid, and mitral areas, among up to nine potential sites. The dataset categorizes these sounds into two primary classes: normal and abnormal. Normal heart sounds were obtained from healthy subjects, while abnormal sounds were recorded from patients diagnosed with cardiac ailments, predominantly heart valve defects such as mitral valve prolapse, mitral regurgitation, aortic stenosis, and post-valvular surgery conditions, as well as coronary artery disease.
The audio recordings, inclusive of contributions from both children and adults, were uniformly truncated to five seconds. Spectrograms of each truncated audio were generated using a window size of 0.061 seconds with a 70% overlap. This multivariate dataset is structured with each dimension representing a frequency band derived from the spectrogram. There are 113 instances in the normal class and 296 in the abnormal class.
B.5 JapaneseVowels
The Japanese Vowels dataset Kudo et al. , sourced from the UCI Machine Learning Repository, comprises recordings from nine male speakers who pronounced the Japanese vowels ‘a’ and ‘e’. Each utterance was analyzed using a 12-degree linear prediction to extract a 12-dimensional time-series representation, with lengths varying originally from 7 to 29. For consistency, all instances in the dataset have been padded to the maximum length of 29. The objective of the classification task is to identify the speaker; hence each 12-by-29 instance matrix is associated with a single class label, ranging from 1 to 9. This dataset serves as a benchmark for assessing the efficacy of time-series classification models in distinguishing speakers based on LPC cepstrum coefficients obtained from their speech patterns.
The dataset includes a total of 640 time-series instances. A training set consists of 30 utterances per speaker, totaling 270 instances. The test set, however, comprises 370 instances and varies in distribution—ranging from 24 to 88 instances per speaker—owing to external factors such as timing and availability during the experimental setup.
B.6 PEMS-SF
The PEMS-SF dataset Cuturi (2011) contains 15 months of daily data sourced from the California Department of Transportation. This dataset details the occupancy rates, ranging from 0 to 1, across various car lanes on the freeways of the San Francisco Bay Area. The data spans from January 1, 2008, to March 30, 2009, with measurements taken every 10 minutes. Each day is treated as an individual time series with a dimension of 963, corresponding to the number of sensors that consistently functioned throughout the observation period. The length of each time series is 144 data points (6 per hour x 24 hours). The dataset excludes public holidays and two anomalous days (March 8, 2009, and March 9, 2008) when sensors recorded no data between 2:00 and 3:00 AM, resulting in a total of 440 valid time series. The classification task involves identifying the day of the week for each series, labeling them with integers from 1 (Monday) to 7 (Sunday). Each attribute within a record reflects the occupancy rate recorded by a sensor at a specific timestamp throughout the day.
B.7 SelfRegulationSCP1
The SelfRegulationSCP1 dataset, sourced from Birbaumer et al. (1999), involves recordings from a healthy subject who was instructed to control a cursor on a screen using cortical potentials. This process was facilitated by tracking the subject’s slow cortical potentials (Cz-Mastoids), where cortical positivity resulted in downward cursor movements and cortical negativity caused it to move upward. Each trial, lasting six seconds, was designed to capture these dynamics, with visual feedback provided between the second 2 and 5.5 of the trial. During each trial, a goal was visually indicated at either the top or bottom of the screen starting from 0.5 seconds to the end of the trial, guiding the subject to generate negative or positive potentials correspondingly. The usable data for each trial, however, spans only 3.5 seconds—from the second 2 to 5.5—corresponding to 896 samples per channel given the sampling rate of 256 Hz.
Data capture involved a PsyLab EEG8 amplifier and a PCIM-DAS1602/16 A/D converter, recording over channels positioned according to the 10/20 system. The dataset includes a training set of 268 trials—168 from the first day and 100 from the second, mixed randomly—and 293 test instances, with class labels indicating positivity or negativity.
B.8 SelfRegulationSCP2
The SelfRegulationSCP2 dataset Birbaumer et al. (1999) includes data from an artificially respirated ALS patient who was tasked with controlling a cursor on a computer screen using cortical potentials. Auditory and visual cues were used to guide the patient, with slow cortical potentials measured at the Cz-Mastoids. A positive potential moved the cursor downward, whereas a negative potential moved it upward. Each trial lasted 8 seconds, with the cursor movement direction (up for negativity, down for positivity) indicated both visually and auditorily from the 0.5 to 7.5 second marks. Auditory instructions were given precisely at the 0.5-second mark, and visual feedback was available from seconds 2 to 6.5. Only the data from this 4.5-second feedback period, translating to 1152 samples per channel at a 256 Hz sampling rate, are used for training and testing.
EEG data were collected from several sites according to the 10/20 system and included channels for detecting vertical eye movements (vEOG). The EEG signals were not corrected for EOG artifacts, providing a raw view of the cortical activity. The dataset comprises 200 trials for training, evenly split between two classes, and an additional 180 trials for testing, recorded on the same day but after the training session data. Each trial spans 7 dimensions and a series length of 1152.
B.9 Spoken Arabic Digits
The Spoken Arabic Digits dataset Bedda & Hammami (2010) consists of 8,800 time series data entries derived from the vocal utterances of 88 native Arabic speakers (44 males and 44 females, aged between 18 and 40). Each dataset entry represents one of ten Arabic digits, spoken ten times by each speaker. The dataset captures 13 Mel Frequency Cepstral Coefficients (MFCCs) for each sound snippet, which are extracted under the following audio processing conditions:
  • Sampling rate: 11025 Hz
  • Bit depth: 16 bits
  • Window function: Hamming
  • Pre-emphasis filter: $1-0.97Z^{-1}$
Each line in the database corresponds to one frame of analysis, listing the 13 MFCCs separated by spaces. These coefficients effectively capture the spectral properties essential for recognizing spoken digits. This structured approach facilitates robust time-series analysis for speech recognition tasks involving Arabic numerals.
B.10 UWaveGestureLibrary
The UWaveGestureLibrary Liu et al. (2009) comprises a set of eight simple gestures, each generated from accelerometer data. The dataset records the X, Y, and Z coordinates corresponding to each gesture’s motion. Every time series within this dataset consists of 315 data points.
Appendix C Model Performance
Figure 3 displays a comparison of models based on the average classification accuracy across all datasets, as summarized in Table 6.
Figure 3: Model comparison based on classification accuracy, averaged across all datasets listed in Table 6. For detailed results, refer to Table 1.
Appendix D Computational Cost Analysis
Table 7 discusses the training and inference times for our model as compared to the SOTA OneFitsAll (Zhou et al., 2024) across the 10 benchmark datasets.
Table 7: Comparison of training and inference times (in seconds) per batch for OneFitsAll (Zhou et al., 2024) versus our model across the 10 benchmark datasets. The Delta (%) is calculated using the formula 100×Cost with Our ModelCost with $100\times\frac{\text{Cost with Our Model}}{\text{Cost with OneFitsAll}}$, quantifying the computational efficiency of our model relative to OneFitsAll.
[…]
Time
Training time ($s$) LETS-C (Ours) 0.01 0.01 0.009 0.01 0.01 0.11 0.004 0.01 0.01 0.01 0.02
OneFitsAll 0.35 0.23 0.14 0.07 0.02 0.28 0.11 0.13 0.97 0.07 0.24
Delta (%) 3.17 8.02 6.71 16.18 42.85 40.49 4.30 8.40 1.21 15.30 14.66
Inference time ($s$) LETS-C (Ours) 0.01 0.05 0.01 0.01 0.01 0.10 0.004 0.01 0.02 0.01 0.02
OneFitsAll 0.16 0.10 0.06 0.04 0.02 0.15 0.06 0.08 0.15 0.04 0.09
Delta (%) 10.58 50.50 14.75 31.15 68.41 66.10 7.02 19.59 18.00 30.55 31.67
We observe that our approach significantly reduces the total training time to just 14.66% and the inference time to 31.67% of those required by OneFitsAll.
Appendix E Assessing Text Embeddings with Cosine Similarity
Figure 4 visualizes the within-class and between-class cosine similarities of text embeddings derived from the testing time series. Each matrix entry is scaled using min-max normalization to range from 0 to 1, where warmer colors in the heatmap represent higher similarities and darker shades indicate lower similarities. Diagonal entries show within-class similarities, highlighting intra-class cohesion, while off-diagonal entries reveal between-class relationships
Figure 4: Heatmaps illustrating within-class and between-class cosine similarities of text embeddings derived from the testing time series data in Japanese Vowels (left) with 9 classes and Spoken Arabic Digits (right) with 10 classes. On both axes, x and y represent different classes. Diagonal entries indicate within-class similarities, and off-diagonal entries represent between-class similarities. Warmer colors signify higher cosine similarities, while cooler colors suggest lower similarities.
Similar to the training set, we observe that within-class similarity consistently exceeds across-class similarity, thereby validating the hypothesis that text embeddings effectively retain and convey significant information from the underlying time series data.
Appendix F Numerical Precision for Tokenization
To explore the impact of numerical precision on the computation of embeddings, we analyzed classification accuracy across precisions 1 to 6 using four datasets: Handwriting, Heartbeat, JapaneseVowels, and UWaveGestureLibrary. We selected these datasets because they were the smaller ones among the 10 available, making the computation of embeddings more computationally affordable. Figure 5 illustrates the average classification accuracy (%) across these numerical precisions.
Figure 5: Average classification accuracy (%) across numerical precisions 1 to 6. Results are averaged from four datasets: Handwriting, Heartbeat, JapaneseVowels, and UWaveGestureLibrary. See Table 8 in the Appendix for detailed results.
Detailed results can be found in Table 8.
Table 8: Numerical Precision for Tokenization: This table reports classification accuracy (%). Red: Best performance.
Dataset / Precision 1 2 3 4 5 6
Handwriting 23.2 15.9 11.6 11.1 12.0 11.2
Heartbeat 78.0 78.5 79.0 78.5 79.5 78.0
JapaneseVowels 99.2 98.4 97.6 96.8 95.9 95.1
UWaveGestureLibrary 90.6 90.6 92.5 89.1 90.3 91.9
Average $\uparrow$ 72.75 70.85 70.175 68.875 69.425 69.05
AvgWins % $\uparrow$ 50% 0% 25% 0% 25% 0%
Studies in the NLP domain have shown that longer inputs do not perform well with language models Press et al. (2020); Levy et al. (2024). Our empirical analysis supports this claim, revealing a decrease in classification accuracy with increased numerical precision when computing language embeddings. The average accuracy starts at 72.8% with precision 1 and declines to 69% at precision 6. Additionally, the percentage of AvgWins is highest at precision 1. As numerical precision increases, so does the length of the time series and the input to the language embedding model. Note that the maximum token length for text-embedding-3-large embeddings is 8191, and thus keeping precision of 1 ensures that context length doesn’t exceed the maximum permissible token length. This issue is especially problematic for datasets with longer time series, such as the EthanolConcentration dataset, which includes 1751 time steps per sample. This finding led us to opt for precision 1 in our study.
Note that these precision results also depend on the type of tokenization selected, and defining an appropriate tokenization for time series is one of the potential future directions for this work.
Appendix G Embedding Type
Table 9 provides a summary of selected embedding models used in the study.
Table 9: Summary of selected embedding models used in the study.
MTEB Rank Model Embedding Dimensions Max Token Length
15 text-embedding-3-large OpenAI (2024) 3072 8191
6 e5-mistral-7b-instruct Wang et al. (2023) 4096 32768
9 gte-large-en-v1.5 Li et al. (2023) 1024 8192
35 nomic-embed-text-v1 Nussbaum et al. (2024) 768 8192
Appendix H Trade-offs: Model Accuracy vs. Parameter Complexity
Table 10 illustrates the trade-off between model accuracy and the complexity of training parameters in our model, which utilizes both embeddings and time series data as inputs to a lightweight framework.
Table 10: Trade-off between model accuracy and the complexity of training parameters. The accuracy and parameters of the best model are highlighted in bold. The accuracy difference is calculated as the raw difference between the accuracies of the reduced model and the best model. The % Delta in accuracy and parameters is defined separately for each as 100×Accuracy of the reduced modelAccuracy of the optimal $100\times\frac{\text{Accuracy of the reduced model}}{\text{Accuracy of the optimal model}}$ and 100×Parameters of the reduced modelParameters of the optimal $100\times\frac{\text{Parameters of the reduced model}}{\text{Parameters of theoptimal model}}$, quantifying the accuracy and computational efficiency of the reduced model relative to our best model.
[…]
EthanolConcentration 52.9 283950 - -
46 105344 -6.9 |||| 86.95 37.09
FaceDetection 68.9 3842 - -
68.6 2402 -0.3 |||| 99.56 62.51
67.9 962 -1.0 |||| 98.54 25.03
66.4 482 -2.5 |||| 96.37 12.54
Handwriting 23.8 154526 - -
23.2 107226 -0.6 |||| 97.47 69.39
22.7 53626 -1.1 |||| 95.37 34.70
20.2 20394 -3.6 |||| 84.87 13.19
19.4 13426 -4.4 |||| 81.51 8.68
Heartbeat 78 46426 - -
77.6 34820 -0.4 |||| 99.48 75.00
75.1 26908 -2.9 |||| 96.28 57.95
Japanese Vowels 99.2 148233 - -
98.9 123401 -0.3 |||| 99.69 83.24
98.6 105353 -0.6 |||| 99.39 71.07
98.4 100857 -0.8 |||| 99.19 68.03
PEMS-SF 93.1 564231 - -
91.3 173866 -1.8 |||| 98.06 30.81
90.8 85210 -2.3 |||| 97.52 15.10
87.9 69077 -5.2 |||| 94.41 12.24
87.3 62901 -5.8 |||| 93.77 11.14
Self-Regulation SCP1 93.2 302626 - -
92.5 99657 -0.7 |||| 99.24 32.93
Self-Regulation SCP2 62.8 334402 - -
58.9 166306 -3.9 |||| 93.78 49.73
57.8 111106 -5.0 |||| 92.03 33.22
57.2 83330 -5.6 |||| 91.08 24.91
56.1 76386 -6.7 |||| 89.33 22.84
Spoken Arabic Digits 99.2 141790 - -
99 70066 -0.2 |||| 99.79 49.41
98.4 46658 -0.8 |||| 99.19 32.90
98.1 30964 -1.1 |||| 98.89 21.83
98 20646 -1.2 |||| 98.79 14.56
97.8 15487 -1.4 |||| 98.58 10.92
97.6 10328 -1.6 |||| 98.38 7.28
96.6 5308 -2.6 |||| 97.37 3.74
UWave Gesture Library 90.6 263338 - -
88.4 211556 -2.2 |||| 97.57 80.33
87.5 176298 -3.1 |||| 96.57 66.94
This trade-off between model accuracy and parameter complexity is data-dependent. However, we generally observe a trend where a reduction in parameters leads to only a slight decrease in accuracy. Next, let’s take a closer look at three datasets: Heartbeat, PEMS-SF, and Spoken Arabic Digits, to understand how these trade-offs manifest in different contexts.
Heartbeat: In the Heartbeat dataset, we retain 99.48% of the optimal model’s accuracy using only 75% of its trainable parameters. Specifically, the optimal model achieves an accuracy of 78% with 46,426 trainable parameters, while the second-best model achieves 77.6% accuracy with 34,820 parameters. Moreover, we retain 96.28% accuracy with a further reduction to 57.95% of the parameters.
PEMS-SF: In the PEMS-SF dataset, the optimal model starts with an accuracy of 93.1% and 564,231 trainable parameters. Reducing the parameters to 173,866 (30.81% of the original), the model maintains 98.06% of its optimal accuracy at 91.3%. Further parameter reductions to 85,210 (15.10%), 69,077 (12.24%), and 62,901 (11.14%) result in accuracies of 90.8%, 87.9%, and 87.3%, respectively. These reductions illustrate that even significant reductions in parameters only lead to a slight decrease in performance.
SpokenArabicDigits: For the Spoken Arabic Digits dataset, reducing the number of trainable parameters generally correlates with a minor decline in accuracy, though the trade-off is modest. The optimal model, achieving an accuracy of 99.2% with 141,790 trainable parameters, shows that even with substantial reductions to 70,066 parameters (49.41% of the original), the accuracy remains high at 99%, retaining 99.79% of the original model’s accuracy. Further reductions to 46,658 (32.90%), 30,964 (21.83%), 20,646 (14.56%), 15,487 (10.92%), 10,328 (7.28%), and 5,308 (3.74%) yield accuracies of 98.4%, 98.1%, 98%, 97.8%, 97.6%, and 96.6% respectively.
These examples highlight that efficiency in terms of trainable parameters does not linearly correspond to a loss in model accuracy across various datasets. While fewer parameters generally lead to a lower accuracy, the decrement is often proportional and manageable, making these models highly suitable for deployment in resource-constrained environments or for applications requiring rapid processing with minimal computational overhead.